# Constrained Portfolio Analysis in High Dimensions: Tracking Error and Weight Constraints

### v1

@@ A.3 Joint Tracking Error and Binding Weight Constraints
To consider the first right side term in (A.73)
$|\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}|/p=O(\bar{s}r_{T}K^{5/2}),$ (A.74)
by Assumption 8, by (B.8) of Caner et al. (2023) $\|\boldsymbol{\mu}\|_{\infty}=O(K)$, and the analysis in (A.59) with $\|\boldsymbol{\mu}\|_{1}/p\leq\|\boldsymbol{\mu}\|_{\infty}$ By assumption $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p|\geq c>0$
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}\right|=O(\bar{s}r_{T}K^{5/2}).$ (A.75)
Then in the same way by using (A.59) and assuming $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p|\geq c>0$
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}\right|=O(\bar{s}r_{T}K^{3/2}).$ (A.76)
@@
$\|\boldsymbol{l}\|_{1}=\|\frac{\boldsymbol{k}-\boldsymbol{a}}{w_{k}-w_{a}}\|_{1}=O(\bar{s}r_{T}K^{3/2}),$
@@
$|\hat{w}_{u}-w_{u}|=\left|\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}\right|+\left|\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}\right|.$ (A.77)
@@
where we use add and subtract $\Theta\hat{\mu}$ and triangle inequality to get the second inequality, and the rates are by (A.20)(A.21) and the (A.20) rate is slower due to $l_{T}$ definition in (2).
Use (A.80) with Assumption $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}|/p\geq c>0$ in (A.79) to have
@@
$|\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p|\left|\frac{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p-\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}{(\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p)(\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p)}\right|=O(\bar{s}r_{T}K^{5/2})O_{p}(\bar{s}l_{T}K)=O_{p}(\bar{s}^{2}r_{T}l_{T}K^{7/2}).$ (A.84)
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}\right|=O_{p}(\bar{s}^{2}r_{T}l_{T}K^{7/2}).$ (A.85)
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}\right|\leq\left|\frac{\boldsymbol{1}_{R}^{\prime}(\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta})^{\prime}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}\boldsymbol{1}_{p}/p}\right|$ (A.86)
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{p}\right|\left|\frac{1}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}-\frac{1}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}\right|=O_{p}(\bar{s}r_{T}K^{3/2})O_{p}(\bar{s}l_{T})=O_{p}(\bar{s}^{2}r_{T}l_{T}K^{3/2}),$ (A.89)
@@
$\|\kappa(\hat{w}_{u}-w_{u})(\hat{\boldsymbol{l}}-\boldsymbol{l})\|_{1}\leq|\kappa||\hat{w}_{u}-w_{u}|\|\hat{\boldsymbol{l}}-\boldsymbol{l}\|_{1}=O_{p}(\bar{s}^{2}r_{T}l_{T}K^{7/2})O_{p}(\bar{s}^{3}r_{T}^{2}l_{T}K^{3}).$ (A.94)
@@
$\|\kappa(\hat{w}_{u}-w_{u})\boldsymbol{l}\|_{1}\leq|\kappa||\hat{w}_{u}-w_{u}|\|\boldsymbol{l}\|_{1}=O_{p}(\bar{s}^{2}r_{T}l_{T}K^{7/2})O(\bar{s}r_{T}K^{3/2})=O_{p}(\bar{s}^{3}r_{T}^{2}l_{T}K^{5}).$ (A.95)
@@
$\|(\omega-\kappa w_{u})(\hat{\boldsymbol{l}}-\boldsymbol{l})\|_{1}\leq[|\omega|+|\kappa||w_{u}|]\|\hat{\boldsymbol{l}}-\boldsymbol{l}\|_{1}=O(\bar{s}r_{T}K^{5/2})O_{p}(\bar{s}^{3}r_{T}^{2}l_{T}K^{3})=O_{p}(\bar{s}^{4}r_{T}^{3}l_{T}K^{11/2}).$ (A.96)
@@
$\|(\omega-\kappa\hat{w}_{u})\hat{\boldsymbol{l}}-(\omega-\kappa w_{u})\boldsymbol{l}\|_{1}=O_{p}(\bar{s}^{4}r_{T}^{3}l_{T}K^{11/2}).$ (A.97)
@@
Proof of Theorem 2(ii). Define $\hat{\boldsymbol{w}}_{R}\mathrel{\mathop{:}}=\hat{\boldsymbol{w}}_{cp}+\boldsymbol{m}_{R}$ and $\boldsymbol{w}_{R}^{*}\mathrel{\mathop{:}}=\boldsymbol{w}_{cp}^{*}+\boldsymbol{m}_{R}$. Then
$\hat{\boldsymbol{w}}_{R}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{R}$displaystyle-- $\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{R}^{*}$
@@
$(\hat{\boldsymbol{w}}_{cp}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{cp})$displaystyle-- $(\boldsymbol{w}_{cp}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{cp}^{*})=(\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*})^{\prime}\boldsymbol{\Sigma}_{y}(\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*})$ (A.99)
@@
$(\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*})^{\prime}\boldsymbol{\Sigma}_{y}(\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*})\leq\|\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*}\|_{1}^{2}\|\boldsymbol{\Sigma}_{y}\|_{\infty}=O_{p}(r_{w2}^{2})O(K^{2}).$ (A.100)
@@
$=$ $O(\bar{s}^{2}r_{T}^{2}K^{4})O(K^{2})O_{p}(r_{w2})=O_{p}(r_{w2}s_{T}^{2}r_{T}^{2}K^{6}),$
where we use the Holder inequality and $\|\boldsymbol{A}\boldsymbol{x}\|_{\infty}\leq\|\boldsymbol{A}\|_{\infty}\|\boldsymbol{x}\|_{1}$ inequality for generic matrix $\boldsymbol{A}$, and generic vector $\boldsymbol{x}$ for the inequality and the rates are from (A.28), and (A.101) with Theorem 2(i). Combine (A.100)(A.102) in (A.99) to have, since $r_{w2}=o(1)$ by Assumption 8
$(\hat{\boldsymbol{w}}_{cp}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{cp})-(\boldsymbol{w}_{cp}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{cp}^{*})=O_{p}(r_{w2}\bar{s}^{2}r_{T}^{2}K^{6}).$ (A.103)
Then given that $\|\boldsymbol{m}_{R}\|_{1}=O(\|\boldsymbol{w}_{cp}^{*}\|_{1})$, and by the Holder inequality and $\|\boldsymbol{A}\boldsymbol{x}\|_{\infty}\leq\|\boldsymbol{A}\|_{\infty}\|\boldsymbol{x}\|_{1}$ inequality for generic matrix $\boldsymbol{A}$, and generic vector $\boldsymbol{x}$ for the inequality, $\|\boldsymbol{m}_{R}\|_{1}\leq\|\boldsymbol{m}\|_{1}$ by $\boldsymbol{m}_{R}$ definition
$|\boldsymbol{m}_{R}^{\prime}\boldsymbol{\Sigma}_{y}(\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp})|$$\leq$ $\|\boldsymbol{m}_{R}\|_{1}\|\boldsymbol{\Sigma}_{y}\|_{\infty}\|\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*}\|_{1}$ (A.104)
$=$ $O(\bar{s}^{2}r_{T}^{2}K^{4})O(K^{2})O_{p}(r_{w2})=O_{p}(r_{w2}\bar{s}^{2}r_{T}^{2}K^{6}),$
@@
$\|\frac{\hat{\boldsymbol{w}}_{R}^{{}^{\prime}}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{R}}{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{R}^{*}}-1\|_{1}=O_{p}(r_{w2}\bar{s}^{2}r_{T}^{2}K^{6})=o_{p}(1).$
@@
$\|\hat{\boldsymbol{w}}_{R}-\boldsymbol{w}_{R}^{*}\|_{1}\|\boldsymbol{\mu}\|_{\infty}=O_{p}(r_{w2}K).$
Also by $\boldsymbol{w}_{R}^{*}\mathrel{\mathop{:}}=\boldsymbol{w}_{cp}^{*}+\boldsymbol{m}_{R}$ and (A.101) and Assumption 8
$\|\boldsymbol{w}_{R}^{*}\|_{1}\|\boldsymbol{\mu}\|_{\infty}=O(\bar{s}^{2}r_{T}^{2}K^{4})O(K)=O(\bar{s}^{2}r_{T}^{2}K^{5}).$
Follow (A.40)-(A.43) analysis, (just replacing $\hat{\boldsymbol{w}}$ with $\hat{\boldsymbol{w}}_{R}$ and $\boldsymbol{w}^{*}$ with $\boldsymbol{w}_{R}^{*}$) with rates in Theorems 2(i)(ii) and (A.101)
$\left|\frac{(\hat{\boldsymbol{w}}^{\prime}_{R}\boldsymbol{\mu})^{2}-(\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\mu})^{2}}{(\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\mu})^{2}}\right|=O_{p}(r_{w2}\bar{s}^{2}r_{T}^{2}K^{5}).$ (A.106)
@@
$\left|\left(\frac{\widehat{SR}_{R}}{SR_{R}}\right)^{2}-1\right|=O_{p}(r_{w2}\bar{s}^{2}r_{T}^{2}K^{6})=o_{p}(1).$

### v2

@@ A.4 Proofs for Joint Tracking Error and Binding Weight Constraints
By the same analysis in (A.20), $r<p$
$|\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}|/p=O(K^{1/2}),$ (A.78)
By Assumption 8, $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p|\geq c>0$
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}\right|=O(K^{1/2}).$ (A.79)
Then in the same way by using (A.63) and assuming $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p|\geq c>0$ as in Assumption 8
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}\right|=O(1).$ (A.80)
@@
$\|\boldsymbol{l}\|_{1}=\|\frac{\boldsymbol{k}-\boldsymbol{a}}{w_{k}-w_{a}}\|_{1}=O(1),$
@@
$|\hat{w}_{u}-w_{u}|\leq\left|\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}\right|+\left|\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}\right|.$ (A.81)
@@
where we use add and subtract $\boldsymbol{\Theta}\hat{\boldsymbol{\mu}}$ and triangle inequality to get the second inequality, and the rates are by (A.24)(A.25) and the (A.24) rate is slower due to $l_{T}$ definition in (2).
Use (A.84) with Assumption 8 $|\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}|/p\geq c>0$ in (A.83) to have
@@
$|\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p|\left|\frac{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p-\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}{(\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p)(\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p)}\right|=O(K^{1/2})O_{p}(\bar{s}l_{T}K)=O_{p}(\bar{s}l_{T}K^{3/2}).$ (A.88)
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\hat{\boldsymbol{\mu}}/p}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{\mu}/p}\right|=O_{p}(\bar{s}l_{T}K^{3/2}).$ (A.89)
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}-\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}\right|\leq\left|\frac{\boldsymbol{1}_{R}^{\prime}(\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta})^{\prime}\boldsymbol{1}_{p}/p}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}\right|$ (A.90)
@@
$\left|\frac{\boldsymbol{1}_{R}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}}{p}\right|\left|\frac{1}{\boldsymbol{1}_{p}^{\prime}\hat{\boldsymbol{\Theta}}^{\prime}\boldsymbol{1}_{p}/p}-\frac{1}{\boldsymbol{1}_{p}^{\prime}\boldsymbol{\Theta}\boldsymbol{1}_{p}/p}\right|=O(1)O_{p}(\bar{s}l_{T})=O_{p}(\bar{s}l_{T}),$ (A.93)
@@
$\|\kappa(\hat{w}_{u}-w_{u})(\hat{\boldsymbol{l}}-\boldsymbol{l})\|_{1}\leq|\kappa||\hat{w}_{u}-w_{u}|\|\hat{\boldsymbol{l}}-\boldsymbol{l}\|_{1}=O_{p}(\bar{s}l_{T}K^{3/2})O_{p}(\bar{s}l_{T}).$ (A.98)
@@
$\|\kappa(\hat{w}_{u}-w_{u})\boldsymbol{l}\|_{1}\leq|\kappa||\hat{w}_{u}-w_{u}|\|\boldsymbol{l}\|_{1}=O_{p}(\bar{s}l_{T}K^{3/2})O(1)=O_{p}(\bar{s}l_{T}K^{3/2}).$ (A.99)
@@
$\|(\omega-\kappa w_{u})(\hat{\boldsymbol{l}}-\boldsymbol{l})\|_{1}\leq[|\omega|+|\kappa||w_{u}|]\|\hat{\boldsymbol{l}}-\boldsymbol{l}\|_{1}=O(K^{1/2})O_{p}(\bar{s}l_{T})=O_{p}(\bar{s}l_{T}K^{1/2}).$ (A.100)
@@
$\|(\omega-\kappa\hat{w}_{u})\hat{\boldsymbol{l}}-(\omega-\kappa w_{u})\boldsymbol{l}\|_{1}=O_{p}(\bar{s}l_{T}K^{3/2}).$ (A.101)
@@
Proof of Theorem 2(ii). Define $\hat{\boldsymbol{w}}_{R}:=\hat{\boldsymbol{w}}_{cp}+\boldsymbol{m}_{R}$ and $\boldsymbol{w}_{R}^{*}:=\boldsymbol{w}_{cp}^{*}+\boldsymbol{m}_{R}$. Then
$\hat{\boldsymbol{w}}_{R}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{R}$ - $\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{R}^{*}$
@@
$(\hat{\boldsymbol{w}}_{cp}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{cp})$ - $(\boldsymbol{w}_{cp}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{cp}^{*})=(\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*})^{\prime}\boldsymbol{\Sigma}_{y}(\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*})$ (A.103)
@@
$(\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*})^{\prime}\boldsymbol{\Sigma}_{y}(\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*})\leq\|\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*}\|_{1}^{2}\|\boldsymbol{\Sigma}_{y}\|_{\infty}=O_{p}(r_{w1}^{2})O(K^{2}).$ (A.104)
@@
$=$ $O(K^{1/2})O(K^{2})O_{p}(r_{w1})=O_{p}(r_{w1}K^{5/2}),$
where we use the Holder inequality and $\|\boldsymbol{A}\boldsymbol{x}\|_{\infty}\leq\|\boldsymbol{A}\|_{\infty}\|\boldsymbol{x}\|_{1}$ inequality for generic matrix $\boldsymbol{A}$, and generic vector $\boldsymbol{x}$ for the inequality and the rates are from (A.32), and (A.105) with Theorem 2(i). Combine (A.104)(A.106) in (A.103) to have, since $r_{w1}=o(1)$ by Assumption 8
$(\hat{\boldsymbol{w}}_{cp}^{\prime}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{cp})-(\boldsymbol{w}_{cp}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{cp}^{*})=O_{p}(r_{w1}K^{5/2}).$ (A.107)
Then given that $\|\boldsymbol{m}_{R}\|_{1}=O(\|\boldsymbol{w}_{cp}^{*}\|_{1})$, and by the Holder inequality and $\|\boldsymbol{A}\boldsymbol{x}\|_{\infty}\leq\|\boldsymbol{A}\|_{\infty}\|\boldsymbol{x}\|_{1}$ inequality for generic matrix $\boldsymbol{A}$, and generic vector $\boldsymbol{x}$ for the inequality,
$|\boldsymbol{m}_{R}^{\prime}\boldsymbol{\Sigma}_{y}(\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*})|$ $\leq$ $\|\boldsymbol{m}_{R}\|_{1}\|\boldsymbol{\Sigma}_{y}\|_{\infty}\|\hat{\boldsymbol{w}}_{cp}-\boldsymbol{w}_{cp}^{*}\|_{1}$ (A.108)
$=$ $O(K^{1/2})O(K^{2})O_{p}(r_{w1})=O_{p}(r_{w1}K^{5/2}),$
@@
$\|\frac{\hat{\boldsymbol{w}}_{R}^{{}^{\prime}}\boldsymbol{\Sigma}_{y}\hat{\boldsymbol{w}}_{R}}{\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\Sigma}_{y}\boldsymbol{w}_{R}^{*}}-1\|_{1}=O_{p}(r_{w1}K^{5/2})=o_{p}(1).$
@@
$\|\hat{\boldsymbol{w}}_{R}-\boldsymbol{w}_{R}^{*}\|_{1}\|\boldsymbol{\mu}\|_{\infty}=O_{p}(r_{w1}K).$
Also by $\boldsymbol{w}_{R}^{*}:=\boldsymbol{w}_{cp}^{*}+\boldsymbol{m}_{R}$ and (A.105) and Assumption 8
$\|\boldsymbol{w}_{R}^{*}\|_{1}\|\boldsymbol{\mu}\|_{\infty}=O(K^{1/2})O(K)=O(K^{3/2}).$
Follow (A.44)-(A.47) analysis, (just replacing $\hat{\boldsymbol{w}}$ with $\hat{\boldsymbol{w}}_{R}$ and $\boldsymbol{w}^{*}$ with $\boldsymbol{w}_{R}^{*}$ with rates in Theorems 2(i)(ii) and (A.105)
$\left|\frac{(\hat{\boldsymbol{w}}^{\prime}_{R}\boldsymbol{\mu})^{2}-(\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\mu})^{2}}{(\boldsymbol{w}_{R}^{*^{\prime}}\boldsymbol{\mu})^{2}}\right|=O_{p}(r_{w1}K^{5/2}).$ (A.110)
@@
$\left|\left(\frac{\widehat{SR}_{R}}{SR_{R}}\right)^{2}-1\right|=O_{p}(r_{w1}K^{5/2})=o_{p}(1).$